OpenStudy (marigirl):

Could someone please help me understand this question and specifically what it means by proportional: Water leaks from the bottom of a tank at a rate proportional to the depth h of the water in the tank. Write down an equation describing this if V is the volume of the water in the tank after time 't' and if the tank is a cylinder show that dh/dt is proportional to h

OpenStudy (anonymous):

do u know anything about Torricelli's law

OpenStudy (marigirl):

if water leaks at the bottom of the tank (dV/dt) at a rate proportional to the depth h of the water in the tank ..how can i write something here..

OpenStudy (anonymous):

\[\frac{ dv }{ dt }=-a \sqrt{2gh}\]

OpenStudy (anonymous):

and

OpenStudy (marigirl):

nope! havent met that before .. what does the right side of the equation mean/stand for?

OpenStudy (anonymous):

\[A(h) \frac{ dh }{ dt}=-k \sqrt{h} \]

OpenStudy (anonymous):

where K is a constant

OpenStudy (marigirl):

Thanks for that

OpenStudy (anonymous):

-a is area btw